# Search for the production of dark gauge bosons in the framework of Einstein-Cartan portal in the simulation of proton-proton collisions at $\sqrt{s}=13.6$ TeV

### v8

@@ Abstract
In the present work, we study the possible production of the heavy neutral dark gauge boson (A) candidates, which originated from a simplified model based on the Einstein-Cartan gravity, in association with dark matter. This study has been performed by studying events with dimuon plus missing transverse momentum produced in the simulated proton-proton collisions at the Large Hadron Collider, at 13.6 TeV center of mass energy and integrated luminosity of 52 fb-1 corresponding to the LHC run 3 circumstances during 2022 and 2023. We provide upper limits, in case no new physics has been discovered, on the masses of various particles in the model as (A), as well as the heavy mediator (torsion field).
@@ I Introduction
The Standard Model of particle physics SM has been tested for more than 40 years [1], and its predictions agree very well with all experimental observations. However, the SM is nowadays considered as a low-energy manifestation of other theories realized at high energy, generically known as BSM (Beyond the Standard Model) theories [2]. One motivation for BSM physics is to have a unified theory for the electromagnetic, weak, and strong interactions, in a unique Grand Unified Theory (GUT) [29]. The Super-Symmetry (SUSY) attempts to also include gravitation lead to models with extra spatial dimensions. These BSM models typically predict the existence of new dark particles at the TeV scale and higher.
The existence of heavy neutral bosons (Z) is a feature of many extensions of the Standard Model. They arise in extended gauge theories, including grand unified theories (GUT) [30], and other models like left-right symmetric models (LRM) [31]. A specific case is the sequential standard model (SSM), in which the Z boson has the same coupling as the SM Z boson [32]. Model of extra dimensions like Randall and Sundrum model (RS) [7] predicts the existence of heavy Kaluza-Klein gravitons. Searches for these heavy dark neutral gauge bosons have been performed at the CMS and ATLAS experiments, at the Large Hadron Collider (LHC), with no evidence of their existence using the full RUN II period of the LHC data taking [33, 34].
@@
Many searches for DM have been performed by analyzing the data collected by the CMS experiment during RUN II. These searches rely on the production of a visible object ”X”, which recoils against the large missing transverse momentum from the dark matter particles leaving a signature of ($\text{X}+E^{\text{miss}}_{T}$) in the detector [18]. The visible particle could be an SM particle like Z [19, 20], W bosons [21, 22] or jets [23, 24], photon [25, 26] or SM Higgs boson [27, 28].
@@
In this analysis, we present a search for dark neutral gauge bosons (A), which originated in a simplified model in the Einstein-Cartan portal, at the LHC simulated proton-proton collisions with 13.6 TeV center of mass energy corresponding to the LHC run 3 circumstances [52]. The topology of the studied simulated events is dimuon, from the decay of A, plus large missing transverse momentum attributed to dark matter. The signal of the above process is very similar to that coming from the decay of Charginos pair into the lepton pair plus MET, carried out by ATLAS [53].
In the following section II, the theoretical formalism of the $U(1)_{D}$ simplified model based on the Einstein-Cartan gravity and its free parameters are presented. Then the simulation techniques used for event generation for the signal and SM background samples are displayed in section III. Afterward, the selection cuts and the strategy of the analysis are explained in section IV. Finally, this analysis’s results and summary are discussed in sections V and VI respectively.
@@ II The simplified model in the framework of Einstein-Cartan gravity
According to the previous equation, the torsion field interacts with both the Standard Model (SM) and hidden fermions through its axial-vector part, with a universal coupling of $\texttt{g}_{\eta}$ = 1/8 at the classical level. This indicates that despite the SM and dark sectors existing in separate realms, the torsion field’s universal coupling to any fermion is essential for generating dark-sector fermions. This occurs when SM fermions collide, leading to the subsequent radiation or decay of lighter dark-sector particles, producing dark gauge bosons with a significantly unsuppressed production cross-section. In the case of the dynamic torsion field, the classical value of $\texttt{g}_{\eta}$ = 1/8 becomes unstable due to quantum corrections, causing $\texttt{g}_{\eta}$ to vary with the energy scale [17]. Due to the torsion mass being in the TeV-scale regime, the $A^{\prime}$ is produced with a high boost, and MET from dark-sector fermions is large where the SM backgrounds are low [17]. For this reason we have chosen $\texttt{g}_{\eta}=0.2$.
@@
where $D_{\mu}=\partial_{\mu}+i\texttt{g}_{\eta}\gamma^{5}S_{\mu}+i\texttt{g}_{D}A^{\prime}_{\mu}$, $M_{\chi}$ is the dark matter mass and $\texttt{g}_{D}$ is the coupling of dark gauge boson to dark matter. The value of the coupling constant $\texttt{g}_{D}$ is taken from [17], which is $\texttt{g}_{D}=1.2$.
@@
$\texttt{g}_{\eta}$ Coupling of torsion field to Dirac fermions       0.2 [17]
$\texttt{g}_{D}$ Coupling of dark gauge boson to dark matter       1.2 [17]
$M_{\chi}$ Mass of dark matter $(\chi)$       $500$ GeV [50, 51]
$M_{A^{\prime}}$ Mass of dark gauge boson $(A^{\prime})$       $[200,...,900]$ GeV
$M_{ST}$ Mass of torsion field $(ST)$       $[1250,...,7000]$ GeV
@@
$M_{A^{\prime}}$ $M_{ST}$ 1250 1500 1750 1800 1970 2000 3000 4000 5000 6000 7000
200 4.3×$\times 10^{-4}$ 86.5×$\times 10^{-4}$ 158.0×$\times 10^{-4}$ 170.0×$\times 10^{-4}$ 180.0×$\times 10^{-4}$ 181.0×$\times 10^{-4}$ 89.5×$\times 10^{-4}$ 26.0×$\times 10^{-4}$ 6.1×$\times 10^{-4}$ 1.2×$\times 10^{-4}$ 2.3×$\times 10^{-5}$
300 9.8×$\times 10^{-7}$ 40.0×$\times 10^{-4}$ 110.0×$\times 10^{-4}$ 120.0×$\times 10^{-4}$ 142.0×$\times 10^{-4}$ 145.0×$\times 10^{-4}$ 87.4×$\times 10^{-4}$ 27.0×$\times 10^{-4}$ 6.6×$\times 10^{-4}$ 1.4×$\times 10^{-4}$ 2.6×$\times 10^{-5}$
400 5.4×$\times 10^{-7}$ 10.0×$\times 10^{-4}$ 65.0×$\times 10^{-4}$ 75.0×$\times 10^{-4}$ 101.0×$\times 10^{-4}$ 105.0×$\times 10^{-4}$ 79.0×$\times 10^{-4}$ 26.0×$\times 10^{-4}$ 6.7×$\times 10^{-4}$ 1.4×$\times 10^{-4}$ 2.8×$\times 10^{-5}$
500 3.3×$\times 10^{-7}$ 8.1×$\times 10^{-7}$ 32.0×$\times 10^{-4}$ $\times 10^{-4}$ 66.0×$\times 10^{-4}$ 69.5×$\times 10^{-4}$ 68.8×$\times 10^{-4}$ 19.0×$\times 10^{-4}$ 6.6×$\times 10^{-4}$ 1.4×$\times 10^{-4}$ 2.8×$\times 10^{-5}$
600 2.1×$\times 10^{-7}$ 3.5×$\times 10^{-7}$ 11.0×$\times 10^{-4}$ 17.0×$\times 10^{-4}$ 39.0×$\times 10^{-4}$ 42.0×$\times 10^{-4}$ 58.0×$\times 10^{-4}$ 23.0×$\times 10^{-4}$ 6.3×$\times 10^{-4}$ 1.4×$\times 10^{-4}$ 2.8×$\times 10^{-5}$
700 1.4×$\times 10^{-7}$ 1.9×$\times 10^{-7}$ 1.2×$\times 10^{-4}$ 4.1×$\times 10^{-4}$ 19.0×$\times 10^{-4}$ 22.0×$\times 10^{-4}$ 48.0×$\times 10^{-4}$ 20.0×$\times 10^{-4}$ 5.9×$\times 10^{-4}$ 1.4×$\times 10^{-4}$ 2.8×$\times 10^{-5}$
800 9.0×$\times 10^{-8}$ 1.2×$\times 10^{-7}$ 1.3×$\times 10^{-7}$ 3.2×$\times 10^{-7}$ 7.1×$\times 10^{-4}$ 9.1×$\times 10^{-4}$ 39.0×$\times 10^{-4}$ 18.0×$\times 10^{-4}$ 5.4×$\times 10^{-4}$ 1.3×$\times 10^{-4}$ 2.7×$\times 10^{-5}$
900 6.0×$\times 10^{-8}$ 7.8×$\times 10^{-8}$ 1.2×$\times 10^{-7}$ 1.4×$\times 10^{-7}$ 1.1×$\times 10^{-4}$ 2.1×$\times 10^{-4}$ 30.0×$\times 10^{-4}$ 16.0×$\times 10^{-4}$ 5.0×$\times 10^{-4}$ 1.2×$\times 10^{-4}$ 2.6×$\times 10^{-5}$
Table 2: The simplified model (based on Einstein Cartan gravity) cross-section measurements times branching ratios (in pb) calculated for different sets of the masses $M_{A^{\prime}}$ (in GeV), and $M_{ST}$ (in GeV), for the mass assumption given in table 1, with dark matter mass ($M_{\chi}=500$ GeV), the following couplings constants $\texttt{g}_{\eta}=0.2,\leavevmode\nobreak\ \texttt{g}_{D}=1.2$ and at $\sqrt{s}=13.6$ TeV.
@@
In this model, there are several free parameters, including the mass of the torsion field ($M_{ST}$), the mass of the dark gauge boson ($M_{A^{\prime}}$), the mass of the dark matter ($M_{\chi}$), and the coupling constants ($\texttt{g}_{\eta}$ and $\texttt{g}_{D}$). As mentioned earlier, the values of these coupling constants are set at $\texttt{g}_{\eta}=0.2$ and $\texttt{g}_{D}=1.2$. By using this combination of coupling values, we can achieve the highest possible cross-section times branching ratios. These model parameters are listed in Table 1.
We are currently investigating the potential production of a heavy neutral dark gauge boson ($A^{\prime}$) at the LHC, with a mass ($M_{A^{\prime}}$) that exceeds 100 GeV. According to the mass requirement specified in table 1, the dark matter mass must be substantial. Previous searches conducted in [50, 51] have indicated that there is a significant chance of detecting dark matter with masses between approximately 100 and 10000 GeV using various search techniques. As a result, we have designated the dark matter mass as $M_{\chi}=500$ GeV.
The typical signature of this process consists of a pair of opposite-sign muons from the decay of A plus a large missing transverse momentum due to the stable dark matter $\chi$. The events we have studied are in the following topology: $\mu^{+}\mu^{-}+E^{miss}_{T}$. This is a clean channel for Standard Model (SM) backgrounds due to the optimization of the CMS detector for this decay channel. Additionally, according to reference [33], the fake rate of the QCD background was much smaller in the dimuon channels compared to the dielectron channel.
@@ III Simulation of signal samples and SM backgrounds
For the simplified model, with the use of mass assumption given in table 1, table 2 indicates the cross-section measurements times branching ratios calculated for different sets of the dark gauge boson (A) and torsion field ($ST$) masses. The simulated signals, used in this analysis, are private production samples, at which we used the matrix element event generator MadGraph5 aMC@NLO v2.6.7 [54].
All the Monte Carlo samples utilized in this analysis and their respective cross sections were computed in next-to-leading order and privately generated. Thus, the contributions of the signal samples and the SM background processes have been estimated from the Monte Carlo simulations, at which they are normalized to their corresponding cross-section and integrated luminosity of 52 fb-1 (the amount of certified data which is collected by CMS experiment so far during LHC run 3) [52].
@@ IV Event selection
The selection of the event, for the analysis, has been designed to reconstruct a final state with two high transverse momentum $(p_{T})$ muons in association with missing transverse momentum, accounting for the DM candidate. The selection is made using cuts applied to different kinematic parameters. Each of the two muons should pass the following pre-selection:
$\bullet$ $p^{\mu}_{T}$ (GeV) $>30$,
$\bullet$ $\eta^{\mu}$ (rad) $<$ 2.4,
[…]
@@
Figure 2: The measured dimuon invariant mass spectrum, for events passing selection 1 summarized in table 3, for the estimated SM backgrounds and different choices of dark gauge boson (A) masses generated based on the simplified model, with mass of torsion field ($M_{ST}=2000$ GeV) and dark matter mass ($M_{\chi}=500$ GeV).
@@
Figure 3: The distribution of the missing transverse momentum, after selection 1; for the expected SM backgrounds, and different A masses produced by the simplified model, with the mass of torsion field ($M_{ST}=2000$ GeV) and dark matter mass ($M_{\chi}=500$ GeV).
@@
Figure 2 shows the distribution of the dimuon invariant mass for events passing selection 1, summarized in table 3; The cyan histogram represents the Drell-Yan background, the yellow histogram represents the vector boson pair backgrounds (WW, WZ, and ZZ), and the red histogram represents the $t\bar{t}$ background. These histograms are stacked. While the signals of the simplified model in the framework of Einstein-Cartan gravity, which have been generated with different masses of the neutral dark gauge boson A with fixed values of the torsion field mass ($M_{ST}=2000$ GeV) and dark matter mass ($M_{\chi}=500$ GeV), are represented by different colored lines, and are overlaid. The corresponding distribution of the missing transverse momentum is presented in Figure 3. It is clearly shown from these figures that the signal samples are overwhelmed by the backgrounds. So, it is necessary to apply a tighter set of cuts to discriminate signals from SM backgrounds, as will be explained in the next paragraph.
@@
In addition to selection 1, extra-tight cuts have been designed to reduce the SM background further. These tight cuts are based on three variables: the first variable is related to the invariant mass of the dimuon, at which we restricted the invariant mass of the dimuon to a small range around the mass of the dark gauge boson A to suppress backgrounds that do not peak at the A mass, a requirement that $250.9\times M_{A^{\prime}}<M_{\mu^{+}\mu^{-}}<M_{A^{\prime}}+$ 25 as suggested in [17]. The second is the relative difference between the transverse momentum of dimuon $(p_{T}^{\mu^{+}\mu^{-}})$ and the missing transverse momentum $(E^{\text{miss}}_{T})$, it has been selected to be less than 0.4. (i.e. $|p_{T}^{\mu^{+}\mu^{-}}-E^{\text{miss}}_{T}|/p_{T}^{\mu^{+}\mu^{-}}<0.4$). The third one is $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}^{\text{miss}}_{T}}$, which is defined as difference in the azimuth angle between the dimuon direction and the missing transverse momentum direction (i.e. $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}^{\text{miss}}_{T}}=|\phi^{\mu^{+}\mu^{-}}-\leavevmode\nobreak\ \phi^{miss}|$ ), it has been selected to be greater than 2.6 rad. These variables effectively reject DY and top quark processes. These adjustments are fine-tuned to maximize the sensitivity of the signal for a range of DM processes.
@@
For dimuon events passing selection 1, we present the distributions of two variables for the signal presentation of the simplified model corresponding to Einstein-Cartan gravity, along with SM backgrounds. The first variable is $|p_{T}^{\mu^{+}\mu^{-}}-E_{T}^{\text{miss}}|/p_{T}^{\mu^{+}\mu^{-}}$, shown in Figure 4(a), while the second variable is $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}_{T}^{\text{miss}}}$, shown in Figure 4(b). The model signals were generated with dark gauge boson mass $M_{A^{\prime}}$ = 200 and 900 GeV, torsion field mass $M_{ST}=2000$ GeV, and dark matter mass $M_{\chi}=500$ GeV. The distributions are scaled to one. The vertical dashed lines in both figures correspond to the chosen cut value for each variable, and these tight cuts have been applied to significantly reduce the SM backgrounds.
@@
IsolationVarRhoCorr $<$ 0.1 IsolationVarRhoCorr $<$ 0.1 IsolationVarRhoCorr $<$ 0.1
@@
$250.9\times M_{A^{\prime}}<M_{\mu^{+}\mu^{-}}<M_{A^{\prime}}+$ 25
@@
Figure 5: The dimuon invariant mass spectrum, for events passing selection 2 listed in table 3, for the estimated SM backgrounds and different choices of dark gauge boson (A) masses generated based on the simplified model, with mass of torsion field ($M_{ST}=2000$ GeV) and dark matter mass ($M_{\chi}=500$ GeV).
The dimuon invariant mass spectrum is shown in Figure 5 for events that meet the criteria listed in Table 3 (selection 2). The histograms represent the estimated SM backgrounds and various dark gauge boson (A) masses generated based on the simplified model, with a torsion field mass of $M_{ST}=2000$ GeV and dark matter mass of $M_{\chi}=500$ GeV.

### v9

@@ Abstract
In this study, we investigate the potential production of a heavy torsion field (TS) at the LHC, which stems from a simplified model rooted in Einstein-Cartan gravity, in connection with dark matter. Within this framework, the torsion field is capable of decaying into pairs of dark matter (DM) particles. Notably, one of these DM particles is heavy enough to decay into dark neutral gauge bosons (A) alongside another DM particle. The Analysis has been performed by studying events with dimuon plus missing transverse energy produced in the simulated proton-proton collisions at the Large Hadron Collider, at 13.6 TeV center of mass energy and integrated luminosity of 52 fb-1 corresponding to the LHC run 3 circumstances during 2022 and 2023. We provide upper limits, in case no new physics has been discovered, on the masses of various particles in the model as (A), as well as the heavy mediator (torsion field).
@@ I Introduction
The Standard Model of particle physics SM has been tested for more than 40 years [1], and its predictions agree very well with all experimental observations. However, the SM is nowadays considered as a low-energy manifestation of other theories realized at high energy, generically known as BSM (Beyond the Standard Model) theories [2]. One motivation for BSM physics is to have a unified theory for the electromagnetic, weak, and strong interactions, in a unique Grand Unified Theory (GUT) [29]. The Super-Symmetry (SUSY) attempts to also include gravitation, leading to models with extra spatial dimensions. These BSM models typically predict the existence of new dark particles at the TeV scale and higher.
The existence of heavy neutral bosons (Z) is a feature of many extensions of the Standard Model. They arise in extended gauge theories, including grand unified theories (GUT) [30], and other models like left-right symmetric models (LRM) [31]. A specific case is the sequential standard model (SSM), in which the Z boson has the same coupling as the SM Z boson [32]. Model of extra dimensions like the Randall and Sundrum model (RS) [7] predicts the existence of heavy Kaluza-Klein gravitons. Searches for these heavy dark neutral gauge bosons have been performed at the CMS and ATLAS experiments, at the Large Hadron Collider (LHC), with no evidence of their existence using the full RUN II period of the LHC data taking [33, 34].
@@
Many searches for DM have been performed by analyzing the data collected by the CMS experiment during RUN II. These searches rely on the production of a visible object ”X”, which recoils against the large missing transverse momentum from the dark matter particles, leaving a signature of ($\text{X}+E^{\text{miss}}_{T}$) in the detector [18]. The visible particle could be an SM particle like Z [19, 20], W bosons [21, 22] or jets [23, 24], photon [25, 26] or SM Higgs boson [27, 28].
@@
In this analysis, we present a search for dark neutral gauge bosons (A), which originated in a simplified model in the Einstein-Cartan portal, at the LHC simulated proton-proton collisions with 13.6 TeV center of mass energy corresponding to the LHC run 3 circumstances [52]. The topology of the studied simulated events is dimuon, from the decay of A, plus large missing transverse energy attributed to dark matter. The signal of the above process is very similar to that coming from the decay of Charginos pair into the lepton pair plus MET, carried out by ATLAS [53].
In the following section II, the theoretical formalism of the $U(1)_{D}$ simplified model based on the Einstein-Cartan gravity and its free parameters are presented. Then the simulation techniques used for event generation for the signal and SM background samples are displayed in section III. Afterward, the selection cuts and the strategy of the analysis are explained in section IV. Finally, this analysis’s results and summary are discussed in sections V and VI, respectively.
@@ II The simplified model in the framework of Einstein-Cartan gravity
According to the previous equation, the torsion field interacts with both the Standard Model (SM) and hidden fermions through its axial-vector part, with a universal coupling of $\texttt{g}_{\eta}$ = 1/8 = 0.125 [17] at the classical level. This indicates that despite the SM and dark sectors existing in separate realms, the torsion field’s universal coupling to any fermion is essential for generating dark-sector fermions. This occurs when SM fermions collide, leading to the subsequent radiation or decay of lighter dark-sector particles, producing dark gauge bosons with a significantly unsuppressed production cross-section.
@@
where $D_{\mu}=\partial_{\mu}+i\texttt{g}_{\eta}\gamma^{5}S_{\mu}+i\texttt{g}_{D}A^{\prime}_{\mu}$, $M_{\chi}$ is the dark matter mass and $\texttt{g}_{D}$ is the coupling of dark gauge boson to dark matter.
@@
$\texttt{g}_{\eta}$ Coupling of torsion field to Dirac fermions
$\texttt{g}_{D}$ Coupling of dark gauge boson to dark matter
$M_{\chi}$ Mass of dark matter $(\chi)$     500500 GeV [50, 51]
$M_{A^{\prime}}$ Mass of dark gauge boson $(A^{\prime})$     [200,,900][200,...,900] GeV
$M_{TS}$ Mass of torsion field $(TS)$     [1250,,7000][1250,...,7000] GeV
@@
$M_{A^{\prime}}$ $M_{TS}$ 1250 1500 1750 1800 1970 2000 3000 4000 5000 6000 7000
@@
[…]
Table 2: The simplified model (based on Einstein Cartan gravity) cross-section measurements times branching ratios (in pb) calculated for different sets of the masses $M_{A^{\prime}}$ (in GeV), and $M_{TS}$ (in GeV), for the mass assumption given in table 1, with dark matter mass ($M_{\chi}=500$ GeV), the following couplings constants $\texttt{g}_{\eta}=0.2,\penalty 10000\ \texttt{g}_{D}=1.2$ and at $\sqrt{s}=13.6$ TeV.
In general, the neutral dark gauge boson (A) subsequently decays to the SM fermion pairs [17]. In this article, the Einstein-Cartan gravity provides a portal for a significant production of $A^{\prime}$ dark gauge boson, which corresponds to a $U(1)_{D}$ symmetry in the dark sector and couples to the SM particles through only the kinetic mixing ($\epsilon\leq 10^{-2}$). For the process $pp\rightarrow TS\rightarrow\chi\bar{\chi}A^{\prime}$, if the decay channels of the $A^{\prime}$ to dark-sector particles are inaccessible, the cross section of the this process depends on the $A^{\prime}$ and torsion masses, the fermion-torsion coupling, the dark gauge boson to dark matter coupling, and the branching ratio Br($A^{\prime}\rightarrow l^{+}l^{-}$), without depending on the kinetic mixing parameter.
[…]
If the torsion mass ($M_{TS}$) is situated at the Planck scale, its interactions would be significantly suppressed, making it nearly undetectable in both current and future collider experiments. For torsion to have a meaningful impact on collider physics, its effective mass must be positioned at energy levels considerably lower than the Planck scale. Therefore, Einstein-Cartan gravity does not inherently necessitate a torsion field with a mass at the Planck scale. Although it is conceptually connected to the Planck scale, the effective scale of torsion is shaped by other elements within the theory, including its coupling to matter [54, 17]. In this model, there are several free parameters, including $M_{TS}$, the mass of the dark gauge boson ($M_{A^{\prime}}$), the mass of the dark matter ($M_{\chi}$), and the coupling constants ($\texttt{g}_{\eta}$ and $\texttt{g}_{D}$). These model parameters are listed in Table 1.
In the context of the dynamic torsion field, the classical value of $\texttt{g}_{\eta}=1/8=0.125$ becomes unstable due to quantum corrections, resulting in fluctuations of $\texttt{g}_{\eta}$ as the energy scale varies [17]. Consequently, we decided to explore several values of $\texttt{g}_{\eta}$, in addition to the classical value of 0.125. Specifically, we included one higher value (0.2) and another lower (0.061 or 0.067).
The chosen benchmark values for the coupling parameter $\texttt{g}_{D}$ are set at 1, in line with the recommendations from the LHC Dark Matter Working Group [55]. Additionally, a value of $\texttt{g}_{D}$ = 1.2 is adopted based on the suggestions put forth by [17].
Since the torsion mass falls within the TeV-scale range, the $A^{\prime}$ is produced with a significant boost, resulting in substantial missing energy (MET) from dark-sector fermions, while the standard model (SM) backgrounds remain low [17].
We are currently investigating the potential production of a heavy neutral dark gauge boson ($A^{\prime}$) at the LHC, with a mass ($M_{A^{\prime}}$) that exceeds 100 GeV. The dark matter mass must be substantial according to the mass requirement specified in table 1. Previous searches conducted in [50, 51] have indicated that there is a significant chance of detecting dark matter with masses between approximately 100 and 10000 GeV using various search techniques. As a result, we have designated the dark matter mass as $M_{\chi}=500$ GeV.
The typical signature of this process consists of a pair of opposite-sign muons from the decay of A plus a large missing transverse energy due to the stable dark matter $\chi$. The events we have studied are in the following topology: $\mu^{+}\mu^{-}+E^{miss}_{T}$. This is a clean channel for Standard Model (SM) backgrounds due to the optimization of the CMS detector for this decay channel. Additionally, according to reference [33], the fake rate of the QCD background was much smaller in the dimuon channels compared to the dielectron channel.
@@ III Simulation of signal samples and SM backgrounds
In the simplified model, we used the mass assumptions outlined in Table 1. In table 2, we present the calculated measurements of cross-sections multiplied by branching ratios for the chosen benchmark values of the coupling parameters, specifically $\texttt{g}_{\eta}=0.2$ and $\texttt{g}_{D}=1.2$, in line with the recommendations from [17]. These calculations for the cross-sections are conducted across various sets of masses for the dark gauge boson (A) and the torsion field ($ST$). The simulated signals, used in this analysis, are private production samples, at which we used the matrix element event generator MadGraph5 aMC@NLO v2.6.7 [56].
All the Monte Carlo samples utilized in this analysis and their respective cross sections were computed in next-to-leading order and privately generated. Thus, the contributions of the signal samples and the SM background processes have been estimated from the Monte Carlo simulations, at which they are normalized to their corresponding cross-section and integrated luminosity of 52 fb-1 (the amount of certified data which is collected by the CMS experiment so far during LHC run 3) [52].
@@ IV Event selection
The selection of the event, for the analysis, has been designed to reconstruct a final state with two high transverse momentum $(p_{T})$ muons in association with missing transverse energy, accounting for the DM candidate. The selection is made using cuts applied to different kinematic parameters. Each of the two muons should pass the following pre-selection:
$\bullet$ $p^{\mu}_{T}$ (GeV) >30>30,
$\bullet$ $\eta^{\mu}$ (rad) << 2.4,
$\bullet$ $\text{IsolationVarRhoCorr}<0.1$,
@@
Figure 2: The measured dimuon invariant mass spectrum, for events passing selection 1 summarized in table 3, for the estimated SM backgrounds and different choices of dark gauge boson (A) masses generated based on the simplified model, with mass of torsion field ($M_{TS}=2000$ GeV) and dark matter mass ($M_{\chi}=500$ GeV).
@@
Figure 3: The distribution of the missing transverse energy, after selection 1; for the expected SM backgrounds, and different A masses produced by the simplified model, with the mass of torsion field ($M_{TS}=2000$ GeV) and dark matter mass ($M_{\chi}=500$ GeV).
@@
Figure 2 shows the distribution of the dimuon invariant mass for events passing selection 1, summarized in table 3; The cyan histogram represents the Drell-Yan background, the yellow histogram represents the vector boson pair backgrounds (WW, WZ, and ZZ), and the red histogram represents the $t\bar{t}$ background. These histograms are stacked. While the signals of the simplified model in the framework of Einstein-Cartan gravity, which have been generated with different masses of the neutral dark gauge boson A with fixed values of the torsion field mass ($M_{TS}=2000$ GeV) and dark matter mass ($M_{\chi}=500$ GeV), are represented by different colored lines, and are overlaid. The corresponding distribution of the missing transverse energy is presented in Figure 3. It is clearly shown from these figures that the signal samples are overwhelmed by the backgrounds. So, it is necessary to apply a tighter set of cuts to discriminate signals from SM backgrounds, as will be explained in the next paragraph.
@@
In addition to selection 1, extra-tight cuts have been designed to reduce the SM background further. These tight cuts are based on three variables: the first variable is related to the invariant mass of the dimuon, at which we restricted the invariant mass of the dimuon to a small range around the mass of the dark gauge boson A to suppress backgrounds that do not peak at the A mass, a requirement that $0.9\times M_{A^{\prime}}<M_{\mu^{+}\mu^{-}}<M_{A^{\prime}}+25$ as suggested in [17]. The second is the relative difference between the transverse momentum of dimuon $(p_{T}^{\mu^{+}\mu^{-}})$ and the missing transverse energy $(E^{\text{miss}}_{T})$, it has been selected to be less than 0.4. (i.e. $|p_{T}^{\mu^{+}\mu^{-}}-E^{\text{miss}}_{T}|/p_{T}^{\mu^{+}\mu^{-}}<0.4$). The third one is $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}^{\text{miss}}_{T}}$, which is defined as difference in the azimuth angle between the dimuon direction and the missing transverse momentum direction (i.e. $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}^{\text{miss}}_{T}}=|\phi^{\mu^{+}\mu^{-}}-\penalty 10000\ \phi^{miss}|$ ), it has been selected to be greater than 2.6 rad. These variables effectively reject DY and top quark processes. These adjustments are fine-tuned to maximize the sensitivity of the signal for a range of DM processes.
@@
For dimuon events passing selection 1, we present the distributions of two variables for the signal presentation of the simplified model corresponding to Einstein-Cartan gravity, along with SM backgrounds. The first variable is $|p_{T}^{\mu^{+}\mu^{-}}-E_{T}^{\text{miss}}|/p_{T}^{\mu^{+}\mu^{-}}$, shown in Figure 4(a), while the second variable is $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}_{T}^{\text{miss}}}$, shown in Figure 4(b). The model signals were generated with dark gauge boson mass $M_{A^{\prime}}$ = 200 and 900 GeV, torsion field mass $M_{TS}=2000$ GeV, and dark matter mass $M_{\chi}=500$ GeV. The distributions are scaled to one. The vertical dashed lines in both figures correspond to the chosen cut value for each variable, and these tight cuts have been applied to reduce the SM backgrounds significantly.
@@
IsolationVarRhoCorr << 0.1 IsolationVarRhoCorr << 0.1 IsolationVarRhoCorr << 0.1
@@
$0.9\times M_{A^{\prime}}<M_{\mu^{+}\mu^{-}}<M_{A^{\prime}}+25$
@@
Figure 5: The dimuon invariant mass spectrum, for events passing selection 2 listed in table 3, for the estimated SM backgrounds and different choices of dark gauge boson (A) masses generated based on the simplified model, with mass of torsion field ($M_{TS}=2000$ GeV) and dark matter mass ($M_{\chi}=500$ GeV).
The dimuon invariant mass spectrum is shown in Figure 5 for events that meet the criteria listed in Table 3 (selection 2). The histograms represent the estimated SM backgrounds and various dark gauge boson (A) masses generated based on the simplified model, with a torsion field mass of $M_{TS}=2000$ GeV and dark matter mass of $M_{\chi}=500$ GeV.